OpenStudy (konradzuse):

find the area of the commmon interior of the polar eq r = 4(1+sin(theta)) and r = 4(1-sin(theta)) Show the integration.

OpenStudy (konradzuse):

0 = 4 - 4sin(theta) 0 = 4+4sin(theta) 0/4 = -4sin(theta) 0/4 = 4sin(theta) 4 = sin(theta) -4 = sin(theta) theta = arcsin(4) theta = arcsin(-4)

OpenStudy (konradzuse):

@sami-21 :)

OpenStudy (konradzuse):

@satellite73 do you know about this? :)

OpenStudy (anonymous):

libniz is answering =P

OpenStudy (konradzuse):

We can hope :).

OpenStudy (anonymous):

i know the general idea of what you need to do

OpenStudy (anonymous):

alright, this is what I am going by http://www.wolframalpha.com/input/?i=polar+%7Br+%3D+4%281%2Bsin%28theta%29%29+and+r+%3D+4%281-sin%28theta%29%29%7D \[\int_{0}^{\pi/2}\int_{0}^{4-4sin(\theta)}rdrd\theta\] this takes care of top half top half \[\int_{\pi/2}^{\pi}\int_{0}^{4+4sin(\theta)}rdrd\theta\]

OpenStudy (konradzuse):

all I need to know is if I need to do 2 integrals or where I go from here...

OpenStudy (anonymous):

first graph to see what the common area of both eqautions are

OpenStudy (anonymous):

and libniz if the areas are the same because it being symetrical you don't have to do ... you can do one

OpenStudy (anonymous):

and multiply by 2

OpenStudy (anonymous):

good idea @Outkast3r09

OpenStudy (konradzuse):

my graph is weird.... It goes to 8, this only goes to 6...?

OpenStudy (anonymous):

so take either integral and double it

OpenStudy (konradzuse):

The last question went to 12 like WTF?

OpenStudy (anonymous):

use agraphing calc

OpenStudy (anonymous):

look at my link

OpenStudy (konradzuse):

I did it goes from 0 to 8.

OpenStudy (anonymous):

0? to 8 it should go from thetas

OpenStudy (konradzuse):

ok, well I'm not sure what to do then..... How do I graph it...?

OpenStudy (anonymous):

r is like the radius of the line to the point

OpenStudy (anonymous):

theta is the angle and t is the time

OpenStudy (konradzuse):

I thought I was supposed to graph what was from wolfram...

OpenStudy (anonymous):

that's the graph

OpenStudy (konradzuse):

or any graphing utility.

OpenStudy (konradzuse):

Yeah, so y = 8 at the top point, what I'm saying is the graph I'm provided goes from -6 to 6 for both x,y.

OpenStudy (konradzuse):

that's what I'm graphing right?

OpenStudy (konradzuse):

The integral you said I can just do 1 right?

OpenStudy (konradzuse):

\[2\int\limits_{0}^{\pi/2} 4\sin(\theta)+4 d \theta\]

OpenStudy (anonymous):

so at the top theta=pi/2 r=4+4 sin(pi/2) r=4+4(1)=8

OpenStudy (konradzuse):

mhmhmhm

OpenStudy (anonymous):

yeah , just do one integral and double the answer

OpenStudy (konradzuse):

I'm going to jujst redo the graph screw it... :)

OpenStudy (konradzuse):

ok so I set the integral up correctly? :)

OpenStudy (anonymous):

\[2\int_{0}^{\pi/2}\int_{0}^{4+4sin(\theta)}rdrd\theta\] take the inner integral integral of r is r^2/2 , plug in limit value

OpenStudy (konradzuse):

I've never seen integrals set up like that...

OpenStudy (anonymous):

that's double integral

OpenStudy (konradzuse):

hmm.....? lol

OpenStudy (konradzuse):

so it's r dr and then that dtheta?

OpenStudy (anonymous):

yep

OpenStudy (konradzuse):

so we found r to be 8 then? from what you said? so we do the integral of 8?

OpenStudy (anonymous):

no, that's just graphing to see the graph;

OpenStudy (konradzuse):

ok haha.

OpenStudy (konradzuse):

of 4(sin(theta)+1)?

OpenStudy (konradzuse):

:(

OpenStudy (anonymous):

yes

OpenStudy (konradzuse):

:)

OpenStudy (konradzuse):

and we do that from 0 to pi/2 right? or 0 to 4+4sin(θ)...?

OpenStudy (anonymous):

you intergrate the inner part first

OpenStudy (konradzuse):

I got 4 θ-4 cos(θ)?

OpenStudy (anonymous):

\[\int_{0}^{4+4sin(\theta)}rdr\] do this first

OpenStudy (konradzuse):

from 0 to 4+4sin(theta) hmm..

OpenStudy (anonymous):

I got something more complicated http://www.wolframalpha.com/input/?i=int%5Br%2C%7Br%2C0%2C4%2B4sin%28theta%29%7D%5D

OpenStudy (konradzuse):

wtf.....

OpenStudy (konradzuse):

loooks better than what I got....

OpenStudy (anonymous):

let me do the whole thing

OpenStudy (konradzuse):

thanks :).